a triangle with 14 cenimeters 48 cenemiters and 50 cenimenters what knid of angles is this triangle
Answer by jojo14344(1512)

It is a right triangle.
Why? Right Triangle has hypotenuse side, opposite side, and an adjacent side right? And it follows the pythagorean theorem, (hyp)^2=(opp)^2+(adj)^2. (noting the hypotenuse is always the longest side).
Substituting the given to the pythagorean theorem we have,
(50)^2=(48)^2 + (14)^2
2500= 2304 + 196
2500=2500
Well you see the given fullfills the Pyth. Theo. that's why it is a right triangle.
now for the inside angles, We say we have ABC right triangle, being angle C the 90 deg. Then AB=50cm (hypotenuse), also BC=48cm and AC=14cm.
So for angle A, sinA= BC(opp)/AB(hyp)=48/50
A=sin^-1 (48/50)
A=73.74 deg
And for angle B, sinB=AC(opp)/AB(hyp)=14/50
B=sin^-1(14/50)
B=16.26 deg
To check, ABC=180; 73.74+16.26+90=180
180=180
Note: IN finding angleB, opposite side changes to AC. You can use BC if you want but you need just need to change the trigo.function to cosine. Try it and it willb the same answer. If you need more info for this email me.
Thank you,
Jojo
jojo14344@hotmail.com